th following reaction: K,-1.8 × 10-s at 25°C Submit Incorrect; Try Again; 5 attempts remaining Prade Fnestoack Continue
Answer #1

HC2H3O2(aq) + H2O(l) ---------> H3O^+ (aq) + C2H3O2^- (aq)
I     0.2                             0             0
C     -x                              +x            +x
E     0.2-x                            +x            +x

      Ka   = [H3O^+][C2H3O2^-]/[HC2H3O2]
      1.8*10^-5 = x*x/0.2-x
      1.8*10^-5*(0.2-x) = x^2
         x = 0.0018
        [H3O^+] = 0.0018M
         PH     = -log[H3O^+]
                = -log0.0018
                = 2.7447